# Bayesian brain mapping: population-informed individualized functional topography and connectivity

### v1

@@ Abstract
The spatial topography of brain functional organization is increasingly recognized to play an important role in cognition and disease. Accounting for individual differences in functional topography is also crucial for accurately distinguishing spatial and temporal aspects of brain organization. Yet, accurate estimation of individual functional brain networks from functional magnetic resonance imaging (fMRI) without extensive scanning remains challenging, due to low signal-to-noise ratio. Here, we describe Bayesian brain mapping (BBM), a technique for individual functional topography and connectivity leveraging population information. Population-derived priors for both spatial topography and functional connectivity based on existing spatial templates, such as parcellations or continuous network maps, are used to guide subject-level estimation and combat noise. BBM is highly flexible, avoiding strong spatial or temporal constraints and allowing for overlap between networks and heterogeneous patterns of engagement. Unlike multi-subject hierarchical models, BBM is designed for single-subject analysis, making it highly computationally efficient and translatable to clinical settings. Here, we describe the BBM model and illustrate the use of the BayesBrainMap R package to construct population-derived priors, fit the model, and perform inference to identify engagements. A demo is provided in an accompanying Github repo. We also share priors derived from the Human Connectome Project database and provide code to support the construction of priors from different data sources, lowering the barrier to adoption of BBM for studies of individual brain organization.
@@ 1 Introduction
The functional organization of the brain is highly individualized, both in terms of the spatial configuration (Gordon et al.,, 2017) and temporal dynamics (Finn et al.,, 2015) of functional brain networks. Individual features of functional topography have been shown to be predictive of various phenotypes, as well as disease severity (Bijsterbosch et al.,, 2018; Kong et al.,, 2019; Bijsterbosch et al.,, 2019; Li et al.,, 2022). Therefore, functional topography is a potentially valuable source of imaging-based biomarkers for disease diagnosis and treatment efficacy.
@@
Extracting functional topography accurately at the individual level has historically relied on collecting hours of resting-state functional magnetic resonance imaging (fMRI) on individuals (Braga and Buckner,, 2017; Gordon et al.,, 2017; Xue et al.,, 2021). These “precision neuroimaging” studies have been crucial in establishing the existence of individual differences in spatial topography. However, translational goals, such as the use of functional topography and connectivity features in biomarkers and for personalized treatment, are contingent upon being able to extract these features accurately using moderate amounts of data feasible to collect in clinical settings.
@@
Another motivation for extracting individual function topography from standard-duration scans is its relevance to measurement of functional connectivity (FC), i.e. the temporal synchrony between regions of the brain. The use of group parcellations or network maps misaligned to the individual’s functional topography can lead to biased FC estimates (Smith et al.,, 2011; Bijsterbosch et al.,, 2018, 2019) due to mixing of signals from distinct but proximal functional areas. Thus, true differences in functional topography can be misinterpreted as differences in FC. A potential ramification is that real relationships between behavioral or phenotypic measures and functional topography may be incorrectly attributed to FC. If prediction is our main objective, this mixing of features may not be highly consequential. But if we aim to understand the brain mechanisms underlying certain behaviors or disease states, rather than simply predict them, it is crucial to disentangle differences in spatial topography from differences in temporal engagement and connectivity (Harrison et al.,, 2020).
@@
To extract individual functional topography from typical duration subject-level fMRI data, hierarchical Bayesian models have proven an effective approach. These models reduce noise by combining information from multiple subjects, while respecting individual differences and ensuring correspondence between individuals. Hierarchical models been successfully applied in the context of parcellation (Kong et al.,, 2019, 2021), probabilistic functional modes (PROFUMO) (Harrison et al.,, 2015; Farahibozorg et al.,, 2021), and independent component analysis (ICA) (Guo and Tang,, 2013; Mejia et al.,, 2020). Importantly, these methods have been shown to perform well based on a modest amount of data per individual, potentially avoiding the need to collect prolonged or repeated sessions of data in individuals. This makes them highly attractive in contexts where extensive subject-level scanning is not feasible, including clinical settings where it presents a financial and physical burden for patients.
@@
In addition to the importance of accounting for individual functional topography, there is growing evidence for the existence and relevance of overlapping functional architecture (Faskowitz et al.,, 2020). This favors ”soft” parcellations allowing for overlap between functional networks over traditional ”hard” parcellations (Bijsterbosch et al.,, 2023; Dadi et al.,, 2020; Li et al.,, 2018). Examples of soft parcellation approaches include spatial ICA (Beckmann and Smith,, 2004), temporal ICA (Smith et al.,, 2012), PROFUMO (Harrison et al.,, 2015), non-negative matrix factorization (Margulies et al.,, 2016), gradients (Margulies et al.,, 2016), and dictionaries of functional modes (Dadi et al.,, 2020). What these methods have in common is the use of certain assumptions or constraints to make estimation feasible, though they differ in the specific choice of constraint. For instance, spatial ICA encourages statistical independence between network maps, leading to relatively little overlap between networks. PROFUMO and temporal ICA allow for greater spatial overlap but encourage temporal independence between networks, resulting in low FC between networks (Pervaiz et al.,, 2020). Relaxing these constraints, while continuing to provide sufficient model guidance for estimation accuracy and alignment of features across individuals, may permit both the temporal and spatial features of functional brain organization to be more fully expressed.
@@
Here, we present Bayesian brain mapping (BBM), a pragmatic and flexible hierarchical Bayesian technique for producing individualized functional brain topographic maps without strong constraints on the spatial or temporal structure of the networks. BBM begins with a template, which can take the form of either a group parcellation or a set of continuous network maps. That template is used as the basis for population-derived priors on spatial topography and FC, which are then used to inform subject-level estimation of those same features. BBM is a generalization of template ICA (Mejia et al.,, 2020), which uses group ICA maps as the template, but BBM allows for parcellations, ICA maps, PROFUMO modes, and other types of network maps as the template.
@@
In this work, we describe BBM, illustrate its use, and provide resources to facilitate its adoption. We have three primary goals. First, we demonstrate the use of the BayesBrainMap R package, which includes functions to establish population-derived priors and for subject-level model fitting to derive individual functional topography and connectivity. Second, we illustrate the flexibility of BBM to work with various choices of template, including parcellations and continuous network maps, and how the priors relax the more constrained templates. Finally, we share high-quality population-derived priors for healthy young adults using data from the Human Connectome Project (HCP) (Van Essen et al.,, 2013). These priors can be directly adopted to apply BBM in studies of healthy young adults; for studies of different populations or to use different templates, our pipeline can be adapted to produce custom priors.
@@ 2 Methods
Figure 1 illustrates the BBM framework. The chosen template is first used to construct population-derived priors for the functional topography of and temporal synchrony between networks. These encode inherent population variability in the boundaries, configuration, and overlap of networks, effectively relaxing the constraints of the template. Training subjects from a representative population are used to derive the priors. This training set can comprise holdout data from the focal study or data from a representative repository such as the HCP or Alzheimer’s Disease Neuroimaging Initiative (Mueller et al.,, 2005). The priors are in turn used to guide and inform estimation in a subject-level Bayesian model, avoiding the need to fit a large multi-subject hierarchical model. This ability to analyze data from a single subject at a time makes BBM highly pragmatic, computationally efficient, parallelizable, and clinically feasible.
@@
The BBM population-derived priors reduce noise while retaining signal, resulting in more reliable individual-level network topography maps and the functional connectivity between them (Mejia et al.,, 2020, 2023, 2025). An important feature of the functional topography priors is that the population variance of each network varies across the brain, generally showing greater inter-individual differences where engagement is high, and greater similarity across subjects in background regions (Figure 1). This allows individual differences to be expressed where they exist, while reducing noise in background regions through shrinkage. Due to the continuous, whole-brain nature of BBM network maps, noise reduction in background regions of one network also indirectly contributes to estimation of signal in the same area of the brain in other networks.
The BBM model includes an optional population-derived prior on the FC between networks to encode population trends and variability in connectivity. We recently developed a novel informative prior for correlation matrices based on Cholesky factorization for this purpose (Mejia et al.,, 2025), which accurately captures population patterns of variability in FC. By contrast, the inverse Wishart distribution, a common choice of parametric prior for FC due to its conjugacy with the multivariate Normal distribution (Honnorat et al.,, 2019; Harrison et al.,, 2020; Mejia et al.,, 2025), has limited ability to model variance patterns. The BBM FC prior shrinks the subject-level FC toward the population mean to alleviate sampling error and within-subject variability, building on a long literature of the benefits of Bayesian shrinkage for FC estimation (Ledoit and Wolf,, 2004; Su et al.,, 2009; Chen et al.,, 2010; Varoquaux et al.,, 2010; Shou et al.,, 2014; Mejia et al.,, 2015; Dai et al.,, 2017; Mejia et al.,, 2018; Rahim et al.,, 2019; Pervaiz et al.,, 2020; Honnorat and Habes,, 2022).
In this section, we describe and illustrate the two main steps in BBM: prior estimation and model fitting. We first present the Bayesian model for individual-level functional topography and connectivity with population priors. We then describe how we derive the priors, illustrating the process using data from the HCP. Our HCP-derived priors are freely available for download, and we share our code so it can be adapted to other contexts. We also illustrate application of BBM to individuals from the HCP. All of the techniques described here are implemented in the BayesBrainMap R package, and code for the analyses presented here are available via Github.
@@ 2.1 Hierarchical Bayesian model
For a given subject and session, let ${y}_{tv}$ be the preprocessed BOLD fMRI time series at location $v$ and time point $t$. The BBM model assumes the BOLD time series can be linearly deconstructed into contributions from a set of networks. The model is similar to ICA and other source separation models, but without strong temporal or spatial constraints.
@@
At level (1) of the model, $s_{qv}$ is the spatial engagement of network $q$ at location $v$, and $a_{tq}$ is the temporal activation of network $q$ at time $t$. The vectors $\mathbf{a}_{t}$ and $\mathbf{s}_{v}$ combine those values across all $Q$ networks. The scale of the residual white noise ${e}_{tv}$ is allowed to vary spatially. In ICA and other blind source separation analyses of fMRI data, it is common to include additional components to capture structured noise from head motion and other sources. In BBM, these are not included in the model but are instead estimated and removed beforehand, as described in Mejia et al., (2020). In addition, we recommend the use of decomposition-based denoising like ICA-FIX (Griffanti et al.,, 2014) to reduce structured noise prior to model fitting.
Level (2) of the model incorporates the population-derived prior on the spatial topography in ${s}_{qv}$, where the population mean $s^{0}_{qv}$ and variance $\sigma^{2}_{qv}$ are considered known via the previously estimated prior, and the deviation terms $\delta_{qv}$ represent individual differences (Mejia et al.,, 2020). Spatial dependencies in $\delta_{qv}$ can be modeled via a multivariate spatial prior for additional accuracy and power, though at a higher computational cost (Mejia et al.,, 2023).
Level (3) of the model incorporates a population-derived prior on the FC between networks and is optional. This is accomplished by assuming a multivariate Normal prior on $\mathbf{a}_{t}$ with mean zero and covariance $\mathbf{G}$, and assuming a population-derived hyperprior on $\mathbf{G}$ (Mejia et al.,, 2025). Note that we constrain each column of the mixing matrix to unit variance (for identifiability, as often required in blind source separation), so $Cov(\mathbf{a}_{t})\equiv Cor(\mathbf{a}_{t})$. Hence, the covariance of $\mathbf{a}_{t}$ represents the between-network FC, as the Pearson correlation between the time courses of each network. We provide two choices for the FC prior: the conjugate Inverse-Wishart distribution, which is used in PROFUMO (Harrison et al.,, 2020), or a novel Cholesky prior we developed for this model that accurately encodes patterns of population variance (Mejia et al.,, 2025). The conjugate prior allows for faster computation, while the Cholesky prior improves performance but is more computationally demanding, since it requires sampling from the prior. However, since the model is fit at the single-subject level, it is still fast compared with multi-subject hierarchical models, even with the Cholesky prior.
@@ 2.2 Population-derived priors
Here, we describe the construction of population-derived priors using training data from the HCP and illustrate the process via the BayesBrainMap R package. All steps are detailed in a demo available at https://github.com/mandymejia/BayesianBrainMapping-priors. The HCP-derived priors themselves are available through Open Science Framework (OSF) as described in the Github repo README. These priors can be directly utilized for application of BBM to analyze healthy young adults from the HCP or other datasets, as described in Section 2.3. We do not advise applying HCP-derived priors directly to study individuals from clearly distinct populations, such as children, the elderly, or patients. However, using the process and scripts provided, this workflow can be replicated to produce priors for other populations.
We estimate priors using several different choices of template, including parcellations and continuous network maps (Table 1). Note that whether the template is a hard parcellation or continuous, the priors and subject-level estimates of functional topography are always continuous. Specifically, we use the 17-network Yeo parcellation (Yeo et al.,, 2011), the Midnight Scan Club (MSC) group parcellation (Gordon et al.,, 2017), HCP-derived group ICA maps with 15 to 50 components (Smith et al.,, 2013), and group-level PROFUMO modes. For each template, we build priors with and without global signal regression (GSR), since whether or not to perform GSR is consequential and remains the subject of debate.
@@
GICA 15 GICA 25 GICA 50 PROFUMO 12 Yeo 17 MSC 15
@@
Before estimating the priors, we first select a high quality, balanced subject sample to ensure reliable and representative priors. Starting from the full HCP sample of $N=1206$ subjects, we apply several filters to obtain the final sample for the priors. First, we exclude subjects with insufficient scan duration (<< 10 min) after dropping the first 15 volumes and excluding volumes with excessive head motion, based on a lagged and filtered version of framewise displacement (FD) appropriate for multiband data (Power et al.,, 2019; Pham et al.,, 2023) with a threshold of 0.5mm. Second, we exclude any remaining related subjects. Finally, we balance sex within each age group. After applying all filters, our final sample contains approximately 350 subjects. Two sessions from each subject were then used to train the population-derived priors. Note that priors can also be trained using a single run from each training subject, as described below.
We now briefly describe the process of deriving the priors. For more details, see the Github demo that accompanies this manuscript. All of the steps for prior parameter estimation are implemented in the BayesBrainMap function estimate_prior(). Prior estimation requires a template (a parcellation or set of continuous network maps) and test-retest or single-session fMRI data from a set of $n$ training subjects. If using single-session data, the scan is split down the middle to produce two pseudo sessions. For each subject and session, dual regression (Beckmann et al.,, 2009) is used to produce a noisy set of time courses and spatial maps. For parcellation templates, we use a modified version of dual regression: note that if network maps were binary and non-overlapping, then the first step of dual regression would produce the mean time series within each parcel as the corresponding time series. If group parcels are misaligned to an individual, this will result in mixing signals from multiple networks. To allow for some misalignment between the subject-specific networks and the parcels, we simply compute the median rather than the mean within each parcel. The second step of dual regression then proceeds as usual, as a multiple linear regression model relating the fMRI time series at each location to the network time series to produce a map of engagement for each network. Those test-retest spatial engagement maps and time courses are then used to estimate the parameters for the topography and FC priors in the BBM model, as we now describe.
@@
where $z_{i}$ represent the true, noise-free spatial engagement, and the residuals $e_{ij}$ are assumed independent of $z_{i}$. Our intended prior variance is the between-subject variance, $\sigma^{2}$, while $\sigma^{2}_{e}$ represents noise levels in the dual regression maps. The variance terms in this model can be estimated using standard variance decomposition techniques or iterative estimation procedures for linear mixed effects models.
@@
Estimation of the FC prior in level (3) of the BBM model is based on the $Q\times Q$ Pearson correlation matrices of the dual regression network time series, which represent the FC between networks. Briefly, two options for the prior are available: the inverse-Wishart, which is designed for covariance matrices and has a single parameter controlling the variance, resulting in limited flexibility to capture population variance patterns, and a novel informative prior for correlation matrices based on Cholesky factorizations, which we developed for this model. We refer the reader to Mejia et al., (2025) for details. Our previous analyses found that both choices of prior are beneficial, but the novel permuted Cholesky prior outperforms the inverse-Wishart at identifying reliable FC patterns in individuals (Mejia et al.,, 2025).
@@
A large training sample from a publicly available resource like the HCP is ideal for prior estimation. However, it is sometimes desirable to build a customized prior for the specific population or context of a given study. In that case, it is possible to use a holdout set which has been done successfully in previous studies (Gaddis et al.,, 2022; Derman et al.,, 2025; Butler et al.,, 2025). This holdout set should be large enough to accurately estimate prior parameters, ideally 100 or more, or at least several dozen. It is also possible to use the same subjects for prior estimation and subsequent subject-level analysis, an approach known as empirical Bayes (EB). While EB comes with a risk of overfitting, it also allows for using all of the available data for training, generally resulting in more accurate priors, which may be considered worth the tradeoff. It is important to visually inspect the prior mean and variance maps to assess the level of noise present in the priors. Besides a large training sample, avoiding over-processing of the training data can also reduce noise levels in the priors. Strategies to avoid over-processing include using lenient or data-driven volume censoring, parsimonious nuisance regression/denoising, and high-pass instead of band-pass filtering.
2.3 Fitting the BBM model to single-subject data

### v2

# Bayesian brain mapping: a population-informed framework for personalized functional network topography and connectivity

@@ Abstract
The spatial topography of functional brain organization is increasingly recognized to play an important role in cognition and disease. Accounting for individual differences in functional topography is also crucial for accurately distinguishing spatial and temporal aspects of functional brain connectivity. Yet, accurate estimation of personalized functional brain networks from functional magnetic resonance imaging (fMRI) without extensive scanning remains challenging due to high noise levels. Here, we describe Bayesian Brain Mapping (BBM), a technique for personalized functional topography and connectivity informed by population information. BBM relies on population-derived priors on both spatial topography of networks and between-network functional connectivity to guide subject-level estimation and combat noise. These priors are based on existing spatial templates, such as parcellations or continuous network maps, providing correspondence to those templates. Yet BBM is highly flexible, avoiding strong spatial or temporal constraints and allowing for overlap between networks and heterogeneous patterns of engagement. BBM is designed for single-subject analysis, making it computationally efficient and translatable to clinical settings. Here, we describe the BBM model and illustrate the use of the BayesBrainMap R package to construct population-derived priors, fit the model, and perform inference to identify engagements. A demo is provided in an accompanying Github repo. We also share priors derived from the Human Connectome Project and provide code to support the construction of priors from different data sources, lowering the barrier to adoption of BBM for studies of individual brain organization.
*Contributed equally
Co-author e-mail addresses: ellynbutler2027@u.northwestern.edu, nebel@kennedykrieger.org, afmejia@iu.edu, dderman@iu.edu, ndasilv@iu.edu
ORCIDs: E.R.B. (0000-0001-6316-6444), A.F.M. (0000-0002-4312-8974), D.D. (0000-0002-6895-9248)
Corresponding author: Diego Derman
E-mail: dderman@iu.edu
@@ 1 Introduction
The functional organization of the brain is highly individualized, both in terms of the spatial configuration (Gordon et al., 2017a, ; Braga and Buckner,, 2017; Cui et al.,, 2020) and temporal dynamics (Finn et al.,, 2015) of functional brain networks. Individual features of functional network topography have been shown to be predictive of a range of individual characteristics and conditions including cognition, disease severity, psychopathology, socio-economic status, and neurodevelopment (Bijsterbosch et al.,, 2018; Kong et al.,, 2019; Bijsterbosch et al.,, 2019; Cui et al.,, 2022; Li et al.,, 2022; Keller et al.,, 2023; Lynch et al.,, 2024; Butler et al.,, 2025; Pang et al.,, 2026; Herzberg et al.,, 2024; Zhao et al.,, 2024). Thus, functional network topography is a potentially valuable source of information derived from neuroimaging data for the assessment of disease progression and evaluating the mechanisms underlying treatment efficacy.
@@
Extracting functional topography accurately at the individual level has historically relied on collecting hours of resting-state functional magnetic resonance imaging (fMRI) from each participant (Braga and Buckner,, 2017; Gordon et al., 2017b, ; Xue et al.,, 2021). These precision neuroimaging studies have been crucial in revealing the existence, nature, and extent of individual differences in functional topography. However, translational goals, including using functional topography and connectivity features as biomarkers and for personalized treatment, depend on extracting these features accurately from moderate amounts of data feasible to collect in clinical settings.
@@
Another motivation for extracting individualized functional topography from standard-duration scans is its relevance to measurement of FC, i.e. the temporal synchrony between regions of the brain. The use of group parcellations or network maps misaligned to the individual’s functional topography can lead to biased FC estimates due to mixing of signals from distinct functional areas (Smith et al.,, 2011; Bijsterbosch et al.,, 2018, 2019; Butler et al.,, 2026). Thus, true differences in functional topography can be misinterpreted as differences in FC. A potential ramification is that real relationships between behavioral or phenotypic measures and functional topography may be incorrectly attributed to FC. If prediction is the main objective, such mixing of underlying features may not be highly consequential. But if we aim to understand the brain mechanisms behind certain behaviors or disease states, rather than simply predict them, it is crucial to disentangle differences in spatial topography from differences in temporal engagement and connectivity (Harrison et al.,, 2020).
@@
To extract individualized or “personalized” functional networks (PFNs) from typical duration subject-level fMRI data, hierarchical Bayesian models have proven an effective approach. These models reduce noise by combining information from multiple subjects, while respecting individual differences and ensuring correspondence between individuals. Hierarchical models have been successfully applied in the context of parcellation (Kong et al.,, 2019, 2021), probabilistic functional modes (PROFUMO) (Harrison et al.,, 2015; Farahibozorg et al.,, 2021), and independent component analysis (ICA) (Guo and Tang,, 2013; Mejia et al.,, 2020). Importantly, these methods have been shown to perform well using a modest amount of data per individual, potentially avoiding the need to collect prolonged or repeated fMRI sessions in individuals. This makes them highly attractive in contexts where extensive subject-level scanning is not feasible, including clinical settings where it presents a financial and physical burden for patients.
@@
In addition to the importance of accounting for individual differences in functional topography, there is growing evidence for the existence and relevance of overlapping functional architecture among brain systems or networks (Power et al.,, 2013; Gratton et al.,, 2018; Faskowitz et al.,, 2020; Hermosillo et al.,, 2024). This favors “soft” parcellations allowing for overlap between functional networks over traditional “hard” parcellations (Li et al.,, 2018; Dadi et al.,, 2020; Bijsterbosch et al.,, 2023). Examples of soft parcellation approaches include spatial ICA (Beckmann and Smith,, 2004), temporal ICA (Smith et al.,, 2012), PROFUMO (Harrison et al.,, 2015), non-negative matrix factorization (Li et al.,, 2017), gradients (Margulies et al.,, 2016), and dictionaries of functional modes (Dadi et al.,, 2020). All of these methods use assumptions or constraints to make estimation feasible, though they differ in the specific choice of constraint. For instance, spatial ICA encourages statistical independence between network maps, leading to relatively little overlap between networks. PROFUMO and temporal ICA allow for greater spatial overlap but encourage temporal independence between networks, resulting in low FC between networks (Pervaiz et al.,, 2020). Relaxing these constraints, while continuing to provide sufficient model guidance for estimation accuracy and alignment of features across individuals, may permit both the temporal and spatial features of functional brain organization to be more fully expressed.
@@
Here, we present Bayesian Brain Mapping (BBM), a pragmatic and flexible hierarchical Bayesian technique for producing PFNs without strong constraints on the spatial or temporal structure of the networks. BBM begins with a template, which can be either a group parcellation or a set of continuous network maps. That template is used as the basis for population-derived priors on spatial topography and FC, which are then used to inform subject-level estimation of those same features. BBM is a generalization of template ICA (Mejia et al.,, 2020), which uses group ICA maps as the template, but BBM allows for parcellations, ICA maps, PROFUMO modes, and other types of network maps as the template.
@@
In this work, we describe BBM, illustrate its use, and provide resources to facilitate its adoption. We have three primary goals. First, we demonstrate the use of the BayesBrainMap R package, which includes functions for establishing population-derived priors and for fitting subject-level models to estimate individual functional topography and connectivity. Second, we illustrate the flexibility of BBM to work with various choices of template, including parcellations and continuous network maps, and how the priors relax the more constrained templates. Finally, we share high-quality priors derived from the Human Connectome Project (HCP) (Van Essen et al.,, 2013) using a variety of templates, along with instructions and code to replicate those priors. The priors can be directly adopted to apply BBM to studies of healthy young adults, or our pipeline can be adapted to produce custom priors based on other templates or populations. A comprehensive demo available on Github accompanies this manuscript and provides further details, examples, and visualizations.
@@ 2 Methods
Figure 1 illustrates the BBM framework. First, the selected template is used to construct population-derived priors for the functional topography of and FC between networks. The network priors encode inherent population variability in the boundaries, configuration, and overlap of networks, effectively relaxing any strong constraints of the template. Second, the priors are used to guide and inform estimation in a subject-level Bayesian model to obtain posterior estimates of individual functional topography and FC, as well as measures of uncertainty that can be used for inference. Note that there is no need to fit a large multi-subject hierarchical model. The ability to analyze data from a single subject at a time makes BBM highly pragmatic, computationally efficient, parallelizable, and feasible in clinical settings.
@@
The BBM population-derived priors reduce noise while retaining signal, resulting in more reliable and accurate personalized functional network maps and the FC between them (Mejia et al.,, 2020, 2023, 2025). An important feature of the functional topography priors is that the population variance of each network varies across the brain, generally showing greater inter-individual differences in regions of high engagement, and greater similarity across subjects in background regions where engagement is low (Figure 1). This allows individual differences to be expressed where they exist, while reducing noise in background regions through shrinkage. Due to the continuous, whole-brain nature of BBM network maps, noise reduction in background regions of one network also indirectly contributes to estimation of signal in the same area of the brain in other networks.
BBM includes an optional population-derived prior on the FC between networks to encode population trends and variability in FC, which shrinks the subject-level FC toward the population mean to alleviate sampling error and stabilize within-subject fluctuations. It is well-established that Bayesian shrinkage is beneficial for reliable FC estimation (Ledoit and Wolf,, 2004; Su et al.,, 2009; Chen et al.,, 2010; Varoquaux et al.,, 2010; Shou et al.,, 2014; Mejia et al.,, 2015; Dai et al.,, 2017; Mejia et al.,, 2018; Rahim et al.,, 2019; Pervaiz et al.,, 2020; Honnorat and Habes,, 2022). However, the inverse-Wishart distribution, a common choice of parametric prior for FC due to its conjugacy with the multivariate Normal distribution (Honnorat et al.,, 2019; Harrison et al.,, 2020; Mejia et al.,, 2025), has limited ability to capture population variance patterns. We recently developed a novel informative prior for correlation matrices based on Cholesky factorization, which accurately captures population variance in FC and outperforms the inverse-Wishart for FC estimation (Mejia et al.,, 2025).
In this section, we describe and illustrate the two main steps in BBM: prior estimation and model fitting. We first present the Bayesian model. We then describe how we derive the priors, illustrating the process using HCP data. Our HCP-derived priors are freely available for download, and we share our code so it can be adapted to other contexts. Finally, we illustrate the application of BBM to several individuals from the HCP. All of the techniques described here are implemented in the BayesBrainMap R package, and code for the analyses presented are available via Github (https://github.com/statMINDlab/BBM-demo).
@@ 2.1 Hierarchical Bayesian model
For a given subject, let ${y}_{tv}$ be the preprocessed BOLD fMRI time series at location $v$ and time point $t$. The BBM model assumes the BOLD time series can be linearly deconstructed into contributions from a set of networks. The model is similar to ICA and other source separation models, but without strong temporal or spatial constraints.
@@
At level (1) of the model, $s_{qv}$ is the spatial engagement of network $q$ at location $v$, and $a_{tq}$ is the temporal activation of network $q$ at time $t$. The vectors $\mathbf{a}_{t}$ and $\mathbf{s}_{v}$ combine those values across all $Q$ networks. The scale of the residual white noise ${e}_{tv}$ is allowed to vary spatially. In ICA and other blind source separation analyses of fMRI data, it is common to include additional components to capture structured noise from head motion and other sources. In BBM, these are not included in the model but are automatically estimated and removed beforehand, as described in Mejia et al., (2020). In addition, we recommend decomposition-based denoising, such as ICA-FIX (Griffanti et al.,, 2014) or ICA-AROMA (Pruim et al.,, 2015), to reduce structured noise prior to model fitting as this has been shown to improve performance across methods (Harrison et al.,, 2020).
Level (2) of the model incorporates the population-derived prior on the spatial topography in ${s}_{qv}$, where the population mean $s^{0}_{qv}$ and variance $\sigma^{2}_{qv}$ are considered known, and the deviation terms $\delta_{qv}$ represent individual differences (Mejia et al.,, 2020). Spatial dependencies in $\delta_{qv}$ can be modeled via a multivariate spatial prior for additional accuracy and power, though at a higher computational cost (Mejia et al.,, 2023).
Level (3) of the model incorporates a population-derived prior on the FC between networks and is optional. This is accomplished by assuming a multivariate Normal prior on $\mathbf{a}_{t}$ with mean zero and covariance $\mathbf{G}$, and assuming a population-derived hyperprior on $\mathbf{G}$ (Mejia et al.,, 2025). Note that we constrain each column of the mixing matrix to unit variance (for identifiability, as often required in blind source separation), so $Cov(\mathbf{a}_{t})\equiv Cor(\mathbf{a}_{t})$. Hence, the covariance of $\mathbf{a}_{t}$ is equal to the Pearson correlation between the time courses of each network and represents the between-network FC. We provide two choices for the FC prior: the conjugate inverse-Wishart distribution, which is also used in PROFUMO (Harrison et al.,, 2020), or a novel Cholesky prior developed for this model to more accurately encode population variance patterns (Mejia et al.,, 2025). The conjugate prior allows for faster computation, while the Cholesky prior improves performance but is more computationally demanding, since it requires sampling from the prior. Even with the Cholesky prior, model fitting is quite feasible compared with large hierarchical models.
@@ 2.2 Population-derived priors
Here, we describe the construction of population-derived priors using training data from the HCP and illustrate the process via the BayesBrainMap R package. All steps are detailed in a demo available at https://github.com/statMINDlab/BBM-demo. The HCP-derived priors themselves are available through Open Science Framework (OSF) as described in the Github repo README. These priors can be directly utilized for application of BBM to analyze healthy young adults. While we do not advise applying HCP-derived priors directly to study individuals from clearly distinct populations, such as children, the elderly, or individuals with known neurological conditions, our workflow can be replicated using the process and scripts provided to produce priors for other populations.
We consider several different choices of template, including parcellations and continuous network maps (Table 1). For each template, we build priors with and without global signal regression (GSR), since the decision to use GSR remains the subject of debate. As templates, we use the 17-network Yeo parcellation (Yeo et al.,, 2011), the Midnight Scan Club (MSC) group parcellation (Gordon et al., 2017b, ), HCP-derived group ICA maps with 15 to 50 components (Smith et al.,, 2013), and group-level PROFUMO modes (Harrison et al.,, 2020). Note that whether the template is a hard parcellation or continuous, the BBM priors and individual functional topography maps are always continuous.
@@
GICA 15 GICA 25 GICA 50 PROFUMO 12 Yeo 17 MSC 17
@@
We first select a high-quality/low-motion, balanced training sample to ensure reliable and representative priors. Starting from the full HCP sample of $N=1206$ subjects, we apply several filters. First, we exclude subjects with insufficient scan duration (<< 10 min) after dropping the first 15 volumes and excluding volumes with excessive head motion, based on a lagged and filtered version of framewise displacement (FD) appropriate for multiband data (Power et al.,, 2019; Pham et al.,, 2023) with a threshold of 0.5mm. Second, we exclude any remaining related subjects, retaining the first eligible subject per family after motion and scan-duration filtering. Finally, we balance sex within each age group by randomly downsampling the overrepresented sex so that males and females are equally represented within each age bin. After applying all filters, our final sample contains 348 subjects. Both runs from the first visit were then used to train the population-derived priors. Note that priors can also be trained using a single run per subject, in which case each run is split down the middle to produce two pseudo runs.
We now briefly describe the process of deriving the priors, all steps of which are implemented in the BayesBrainMap function estimate_prior(). For more details, see the Github demo that accompanies this manuscript. For each training subject and session, dual regression (Beckmann et al.,, 2009) is used with the chosen template to produce a noisy set of time courses and spatial maps. [^1] For templates that are parcellations, we use a modified version of dual regression: note that if network maps were binary and non-overlapping, then the first step of dual regression would produce the mean time series within each parcel as the corresponding time series. If group parcels are misaligned to an individual, this will result in mixing signals from multiple networks. To allow for some misalignment between the subject-specific networks and the parcels, we simply compute the median rather than the mean within each parcel. The second step of dual regression then proceeds as usual, as a multiple linear regression modeling the fMRI time series at each location as a function of the network time series to produce a map of engagement for each network. Those test-retest spatial engagement maps and time courses are then used to estimate the parameters for the topography and FC priors in the BBM model, as we now describe.
@@
where $z_{i}$ represent the true, noise-free spatial engagement, and the residuals $e_{ij}$ are assumed independent of $z_{i}$. Our intended prior variance is the between-subject variance of the underlying engagements, $\sigma^{2}$, while $\sigma^{2}_{e}$ represents noise levels in the dual regression maps. The variance terms in this model can be estimated using standard variance decomposition techniques or iterative estimation procedures for linear mixed effects models.
@@
Estimation of the FC prior in level (3) of the BBM model is based on the $Q\times Q$ Pearson correlation matrices of the dual regression network time series. As described above, two options for the prior are available: the inverse-Wishart, which is conjugate but has limited flexibility to capture population variance patterns, and a novel informative prior for correlation matrices based on Cholesky factorizations, which we recently developed. See Mejia et al., (2025) for details of the estimation procedure for both priors.
@@
A large training sample from a publicly available resource like the HCP is ideal for prior estimation, as long as the focal subject or study to be analyzed comes from a similar population. However, it is sometimes desirable to build a customized prior for the specific population or context of a given study. In that case, it is possible to use a holdout set which has been done successfully in previous studies (Gaddis et al.,, 2022; Derman et al.,, 2025). This holdout set should be large enough to accurately estimate prior parameters, ideally 100 or more, or at least several dozen. It is also possible to use the same subjects for prior estimation and subsequent subject-level analysis (Butler et al.,, 2025), an approach known as empirical Bayes (EB). While EB comes with a risk of overfitting, it also allows for using all of the available data for training, generally resulting in more accurate priors, which may be considered worth the tradeoff.
It is vital to visually inspect the prior mean and variance maps to assess the level of noise present. Besides a large training sample, avoiding over-processing of the training data can also reduce noise levels in the priors. Strategies to avoid over-processing include using lenient or data-driven volume censoring, parsimonious nuisance regression/denoising, and high-pass instead of band-pass filtering (Pham et al.,, 2023; Bright et al.,, 2017; Tong et al.,, 2019).
2.3 Fitting the single-subject model